# Decision Mamba: A Multi-Grained State Space Model with Self-Evolution Regularization for Offline RL

### v1

@@ Abstract
While the conditional sequence modeling with the transformer architecture has demonstrated its effectiveness in dealing with offline reinforcement learning (RL) tasks, it is struggle to handle out-of-distribution states and actions. Existing work attempts to address this issue by data augmentation with the learned policy or adding extra constraints with the value-based RL algorithm. However, these studies still fail to overcome the following challenges: (1) insufficiently utilizing the historical temporal information among inter-steps, (2) overlooking the local intra-step relationships among states, actions and return-to-gos (RTGs), (3) overfitting suboptimal trajectories with noisy labels. To address these challenges, we propose Decision Mamba (DM), a novel multi-grained state space model (SSM) with a self-evolving policy learning strategy. DM explicitly models the historical hidden state to extract the temporal information by using the mamba architecture. To capture the relationship among state-action-RTG triplets, a fine-grained SSM module is designed and integrated into the original coarse-grained SSM in mamba, resulting in a novel mamba architecture tailored for offline RL. Finally, to mitigate the overfitting issue on noisy trajectories, a self-evolving policy is proposed by using progressive regularization. The policy evolves by using its own past knowledge to refine the suboptimal actions, thus enhancing its robustness on noisy demonstrations. Extensive experiments on various tasks show that DM outperforms other baselines substantially.
@@ 1 Introduction
Offline Reinforcement Learning (RL) [13, 27, 29, 36] has attracted great attention due to its remarkable successes in the fields of robotic control [5, 34] and games [3, 48]. As transformer [47] has exhibited powerful sequential modeling abilities in natural language processing [4, 41] and computer vision [10, 40], many efforts [6, 8, 59, 25] have been made on applying this architecture to offline RL tasks. Transformer-based methods view the state, action, and reward/return-to-go (RTG) as a sequence, and then predict actions by using the transformer encoder. However, it often fails to make correct decisions when encountering out-of-distribution states or actions, showing limited robustness. Previous work attempts to address this issue from the perspective of data augmentation [49, 62] and objective constraints [6, 51, 59]. However, they introduce a significant number of noises or the overestimation bias. Thus, how to enhance model robustness remains a highly challenging and insufficiently explored issue.
@@ 2.2 State Space Models for Linear-time Sequence Modeling
Recently, State Space Models (SSMs) show high potentials in various domains, including natural language processing [17, 19, 16, 20, 44, 15, 38], computer vision [31, 63, 30, 39, 61] and time-series forecasting [53]. Stemming from signal processing, SSMs capture global dependencies from a sequence more effective in a lightweight structure and shows advantages in compressing the historical information, compared with the transformer architecture. Although SSMs have considerable benefits, it still struggles to perform contextual reasoning. Mamba [15] is thus proposed to alleviate this problem. It introduced a time-varying selective mechanism and a hardware-friendly design, making it as a competitive architecture against with transformer. The Mamba architecture is then adapted to different downstream tasks by considering the characteristics of these tasks. VIM [63] and VMamba [31] introduced 2D SSMs for image understanding. VideoMamba [30] introduced spatio-temporal scan for video understanding. In this work, we take the fine-grained trajectory patterns into consideration, and introduce a multi-grained mamba architecture tailored for RL tasks.
@@ 3.1 Preliminaries
3.2 Decision Mamba
3.2.1 Multi-Grained Mamba
Trajectory Embeddings
$e^{\mathrm{s}}_{i}=\operatorname{MLP}(s_{i}),$ (8)
$e^{\mathrm{R}}_{i}=\operatorname{MLP}(R_{i}),$ (9)
$e^{\mathrm{a}}_{i}=\operatorname{MLP}(a_{i}),$ (10)
Coarse-Grained SSM
[…]
@@ 3.2.1 Multi-Grained Mamba
Further, to better discern the dependencies among SARs within each step, we gather the feature of each single step to obtain the fine-grained representation via a 1D-convolution layer, and then introduce a fine-grained SSM module for extracting the local pattern among SAR, as shown in the middle part and right part of Figure 1. It can be formulated as follows:
@@ 3.2.2 Progressive Self-Evolution Regularization
We adopt the MSE loss, and then the objective 23 is converted to:
@@ 4.2 Overall Results
In order to evaluate our method more comprehensively, we also conduct experiments on other datasets. We adopt representative baselines including BC, DT and LaMo, where LaMo leverages extensive additional natural language corpora and knowledge to enhance the model performance. As illustrated in Table 2, all methods perform exceptionally well on the Expert dataset. However, when it comes to mixing the suboptimal data into the training set, compared with DT, both LaMo and DM exhibit significant superiority, while DM shows a more pronounced overall enhancement. For instance, DM outperforms LaMo by approximately 10% on the Ant-M dataset and around 6% on average. For AntMaze, it requires composing parts of suboptimal trajectories to form more optimal policies for reaching goals. “U-D” is more difficult than “U”, and DM shows superiority in these tasks. More comparison results can be found in Appendix C
@@ 4.3 Ablation Study
To investigate the effectiveness of each component in DM, we conduct experiments with different variants of DM. In particular, we compare 3 different implementations: (1) w/o GB removes the group branch, directly using the original sequence branch in mamba; (2) w/o PSER removes the progressive self-evolution regularization, training the model with the labels in the training dataset; (3) w/o ILO removes the inverse learning objective, only predicting the action in the training procedure. As shown in Table 3, the performance of DM drops significantly without either of these components. Notably, the most substantial performance degradation with about 6% occurs when the PSER module is removed, especially in the suboptimal datasets. This observation verifies the effectiveness of this module in preventing policy from overfitting and thus enhancing its robustness. GB and ILO are also critical for offline RL tasks. Once these two modules are excluded, there is a noticeable reduction in the model’s performance.
@@
Table 3: Ablation Results. “w/o GB/PSER/ILO” represents removing the module of the group branch, the progressive self-evolution regularization, and inverse learning objectives, respectively. Best results are marked in bold.
@@ A.2 Implement Details
For all our experiments, we utilized the default hyperparameter settings and conducted 100,000 training iterations or gradient steps. We implement our method with the official repository of the huggingface.The shared hyperparameters are set to the same as those of LaMo, including the batch size, learning rate, overall training steps, and weight decay. The setting of other specific hyperparameters, including $\beta_{K}$, $\beta_{\mathrm{min}}$ are presented in Table 5. The experiments are conducted on an 8*4090-24G platform, and we run each experiment with four different seeds to ensure its reliability. Each experiment cost around 13 hours with a 4090-24G.
Dataset Learning Rate Weight Decay Context Length Return-to-go Training Steps
$\mathrm{Halfcheetah}$ 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $1800,3600$ 100K
$\mathrm{Hopper}$ 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $8000,12000$ 100K
$\mathrm{Walker}$ 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $2500,5000$ 100K
$\mathrm{Ant}$ 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $3600,6000$ 100K
$\mathrm{Antmaze}$ 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $5,20$ 100K
Table 5: Task-specific Hyperparameters

### v2

@@ Abstract
While the conditional sequence modeling with the transformer architecture has demonstrated its effectiveness in dealing with offline reinforcement learning (RL) tasks, it is struggle to handle out-of-distribution states and actions. Existing work attempts to address this issue by data augmentation with the learned policy or adding extra constraints with the value-based RL algorithm. However, these studies still fail to overcome the following challenges: (1) insufficiently utilizing the historical temporal information among inter-steps, (2) overlooking the local intra-step relationships among states, actions and return-to-gos (RTGs), (3) overfitting suboptimal trajectories with noisy labels. To address these challenges, we propose Decision Mamba (DM), a novel multi-grained state space model (SSM) with a self-evolving policy learning strategy. DM explicitly models the historical hidden state to extract the temporal information by using the mamba architecture. To capture the relationship among state-action-RTG triplets, a fine-grained SSM module is designed and integrated into the original coarse-grained SSM in mamba, resulting in a novel mamba architecture tailored for offline RL. Finally, to mitigate the overfitting issue on noisy trajectories, a self-evolving policy is proposed by using progressive regularization. The policy evolves by using its own past knowledge to refine the suboptimal actions, thus enhancing its robustness on noisy demonstrations. Extensive experiments on various tasks show that DM outperforms other baselines substantially. Our code is available at  https://github.com/aopolin-lv/DecisionMamba.
@@ 1 Introduction
Offline Reinforcement Learning (RL) [13, 27, 29, 38] has attracted great attention due to its remarkable successes in the fields of robotic control [5, 36] and games [3, 32, 50]. As transformer [49] has exhibited powerful sequential modeling abilities in natural language processing [4, 43] and computer vision [10, 42], many efforts [6, 8, 25, 61] have been made on applying this architecture to offline RL tasks. Transformer-based methods view the state, action, and reward/return-to-go (RTG) as a sequence, and then predict actions by using the transformer encoder. However, it often fails to make correct decisions when encountering out-of-distribution states or actions, showing limited robustness. Previous work attempts to address this issue from the perspective of data augmentation [51, 64] and objective constraints [6, 53, 61]. However, they introduce a significant number of noises or the overestimation bias. Thus, how to enhance model robustness remains a highly challenging and insufficiently explored issue.
@@ 2.2 State Space Models for Linear-time Sequence Modeling
Recently, State Space Models (SSMs) show high potentials in various domains, including natural language processing [15, 16, 17, 19, 20, 40, 46], computer vision [30, 31, 33, 41, 65, 63] and time-series forecasting [55]. Stemming from signal processing, SSMs capture global dependencies from a sequence more effective in a lightweight structure and shows advantages in compressing the historical information, compared with the transformer architecture. Although SSMs have considerable benefits, it still struggles to perform contextual reasoning. Mamba [15] is thus proposed to alleviate this problem. It introduced a time-varying selective mechanism and a hardware-friendly design, making it as a competitive architecture against with transformer. The Mamba architecture is then adapted to different downstream tasks by considering the characteristics of these tasks. VIM [65] and VMamba [33] introduced 2D SSMs for image understanding. VideoMamba [30] introduced spatio-temporal scan for video understanding. In this work, we take the fine-grained trajectory patterns into consideration, and introduce a multi-grained mamba architecture tailored for RL tasks.
@@ 3.1 Preliminaries
3.2 Decision Mamba
3.2.1 Multi-Grained Mamba
Trajectory Embeddings
$e^{\mathrm{s}}_{i}=\operatorname{MLP}(s_{i}),\ \ \ \ \ \ e^{\mathrm{R}}_{i}=\operatorname{MLP}(R_{i}),\ \ \ \ \ \ e^{\mathrm{a}}_{i}=\operatorname{MLP}(a_{i}),$ (8)
Coarse-Grained SSM
[…]
@@ 3.2.1 Multi-Grained Mamba
Further, to better discern the dependencies among SARs within each step, we gather the feature of each single step to obtain the fine-grained representation via a 1D-convolution layer, and then introduce a fine-grained SSM module for extracting the local pattern among SAR, as shown in the middle part and right part of Figure 1.
It can be formulated as follows:
@@ 3.2.2 Progressive Self-Evolution Regularization
We adopt the MSE loss, and then the objective 21 is converted to:
@@ 4.2 Overall Results
In order to evaluate our method more comprehensively, we also conduct experiments on other datasets. We adopt representative baselines including BC, DT and LaMo, where LaMo leverages extensive additional natural language corpora and knowledge to enhance the model performance. As illustrated in Table 2, all methods perform exceptionally well on the Expert dataset. However, when it comes to mixing the suboptimal data into the training set, compared with DT, both LaMo and DM exhibit significant superiority, while DM shows a more pronounced overall enhancement. For instance, DM outperforms LaMo by approximately 10% on the Ant-M dataset and around 6% on average. For AntMaze, it requires composing parts of suboptimal trajectories to form more optimal policies for reaching goals. “U-D” is more difficult than “U”, and DM shows superiority in these tasks. More comparison results can be found in Appendix C.
@@ 4.3 Ablation Study
To investigate the effectiveness of each component in DM, we conduct experiments with different variants of DM. In particular, we compare 3 different implementations: (1) w/o MG removes the multi-grained branch, directly using the sequence feature original extracted from mamba; (2) w/o PSER removes the progressive self-evolution regularization, training the model with the labels in the training dataset; (3) w/o ILO removes the inverse learning objective, only predicting the action in the training procedure. As shown in Table 3, the performance of DM drops significantly without either of these components. Notably, the most substantial performance degradation with about 6% occurs when the PSER module is removed, especially in the suboptimal datasets. This observation verifies the effectiveness of this module in preventing policy from overfitting and thus enhancing its robustness. MG and ILO are also critical for offline RL tasks. Once these two modules are excluded, there is a noticeable reduction in the model’s performance.
@@
Table 3: Ablation Results. “w/o MG/PSER/ILO” represents removing the module of multi-grained feature extraction, the progressive self-evolution regularization, and inverse learning objectives, respectively. Best results are marked in bold.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
We would like to thank the reviewers for their constructive comments. This work is supported by Shenzhen College Stability Support Plan (Grant No.GXWD20220817144428005) and National Natural Science Foundation of China (Grant No. 62236003). Additionally, it is also supported by National Natural Science Foundation of China (Grant No. 62406092), and partially supported by Research on Efficient Exploration and Self-Evolution of APP Agents & Embodied Intelligent Cerebellum Control Model and Collaborative Feedback Training Project (Grant No. TC20240403047).
@@ A.2 Implement Details
For all our experiments, we utilized the default hyperparameter settings and conducted 100,000 training iterations or gradient steps. We implement our method with the official repository of the huggingface.The shared hyperparameters are set to the same as those of LaMo, including the batch size, learning rate, overall training steps, and weight decay. The setting of other specific hyperparameters, including $\beta_{K}$, $\beta_{\mathrm{min}}$ are presented in Table 5. The experiments are conducted on an 8*4090-24G platform, and we run each experiment with four different seeds to ensure its reliability.
Dataset Learning Rate Weight Decay Context Length Return-to-go Training Steps $\beta_{K}$ $\beta_{\mathrm{min}}$
Halfcheetah 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $1800,3600$ 100K $0.85$ $0.5$
Hopper 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $8000,12000$ 100K $0.90$ $0.5$
Walker 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $2500,5000$ 100K $0.95$ $0.5$
Ant 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $3600,6000$ 100K $0.85$ $0.5$
Antmaze 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $5,20$ 100K $0.95$ $0.85$
Table 5: Task-specific Hyperparameters.
@@ Appendix E Visualization of Action Distribution
Appendix F Impact/Safegard Statements
Impact Statement
In this study, we propose DM, which effectively extracts historical information, fuses multi-grained information to predict action, and enhances the effectiveness of conditional sequence modeling for offline RL tasks. In addition, we introduce a self-evolving policy learning strategy to effectively prevent the policy from overfitting the noisy trajectories, and further enhance the robustness of the policy. To this end, this technology is expected to advance the offline RL agent which can assist human beings in working under the dangerous circumstances. There are many potential societal consequences of developing advanced RL algorithms, none which we feel must be specifically highlighted here.
Safegrad Statement
In the paper, we have taken rigorous steps to ensure the responsible release of our new offline RL algorithm and any associated models or data. Given the potential for misuse or dual-use of such technology, we have implemented several safeguards to mitigate these risks. Use of our algorithms is subject to the CC BY-NC-SA 4.0 agreement. In addition, we are committed to ongoing monitoring and evaluation of our models’ usage to identify any emerging risks or patterns of misuse. We will take prompt action if we detect any unauthorized or inappropriate use of our technology. Finally, we are open to working with regulators, researchers, and industry partners to further refine our safeguards and ensure the safe and ethical use of our offline RL algorithm.